BioTech Research Center is working to develop a new vaccine for the West Nile Virus. The project is so important that the firm has created three teams of experts to work on the project from different perspectives. Team 1 has a 72 percent chance of success, team 2 an 83 percent chance of success, and team 3 a 80 percent chance. What is the probability that BioTech will develop the vaccine? (Round answer to 4 decimal places, e.g. 15.2500.)
Answer #1

Note: Please rate the answer (upvote) and do ask if any doubt is there.

Answer:

  • The probability that team 1 will fail = 1-0.72= 0.28
  • The probability that team 2 will fail= 1-0.83= 0.17
  • The probability that team 3 will fail= 1-0.8= 0.2
  • The probability that all of them will fail is 0.28*0.17*0.2 = 0.00952
  • This means in case any one of them succeeds we should be able to develop the vaccine. That is given by 1-0.00952= 0.99048
  • The probability of developing vaccine is 0.9905 or 99.0480%

Note: Please rate the answer (upvote) and do ask if any doubt is there.